The RDA of calcium for an adult is 1200 mg/day. Every milligram of a calcium carbonate tablet provides 0.4 mg of elemental calcium. How many 650-mg calcium carbonate tablets are needed to provide the RDA of calcium? Round off the answer upward to the nearest whole number.
Answer #1

Ans) The RDA of calcium for an adult is 1200 mg/day. Every milligram of a calcium carbonate tablet provides 0.4 mg of elemental calcium. 650-mg calcium carbonate tablets are needed to provide the RDA of calcium.

650 × 0.4 = 260 mg

1200 mg= RDA

260 mg = 1

x= 1200/260 = 4.61 ~ 4.6 tablets